OpenStudy (anonymous):

If k(x) = x^2 + 5 and g(x) = sqrt(x-3), what is (k-g)(5)

hero (hero):

1. First understand that \((k - g)(x) = k(x) - g(x)\) 2. Next express \(k(x) - g(x)\): \(k(x) - g(x) = x^2 - 5 - \sqrt{x - 3}\) 3. Finally write the expression for \((k - g)(5)\) \((k - g)(5) = 5^2 - 5 - \sqrt{5 - 3} = 25 - 5 - \sqrt{2} = 20 - \sqrt{2}\)